# Spatially-guided Temporal Aggregation for Robust Event-RGB Optical Flow Estimation

### v1

@@ Abstract
Current optical flow methods exploit the stable appearance of frame (or RGB) data to establish robust correspondences across time. Event cameras, on the other hand, provide high-temporal-resolution motion cues and excel in challenging scenarios. These complementary characteristics underscore the potential of integrating frame and event data for optical flow estimation. However, most cross-modal approaches fail to fully utilize the complementary advantages, relying instead on simply stacking information. This study introduces a novel approach that uses a spatially dense modality to guide the aggregation of the temporally dense event modality, achieving effective cross-modal fusion. Specifically, we propose an event-enhanced frame representation that preserves the rich texture of frames and the basic structure of events. We use the enhanced representation as the guiding modality and employ events to capture temporally dense motion information. The robust motion features derived from the guiding modality direct the aggregation of motion information from events. To further enhance fusion, we propose a transformer-based module that complements sparse event motion features with spatially rich frame information and enhances global information propagation. Additionally, a mix-fusion encoder is designed to extract comprehensive spatiotemporal contextual features from both modalities. Extensive experiments on the MVSEC and DSEC-Flow datasets demonstrate the effectiveness of our framework. Leveraging the complementary strengths of frames and events, our method achieves leading performance on the DSEC-Flow dataset. Compared to the event-only model, frame guidance improves accuracy by 10%. Furthermore, it outperforms the state-of-the-art fusion-based method with a 4% accuracy gain and a 45% reduction in inference time.
@@ I Introduction
Optical flow estimation plays a vital role in understanding object motion between image pairs, providing valuable insights into scene dynamics. Frame-based optical flow methods have leveraged the consistent visual appearance of images, leading to extensive development. Over the past decade, learning-based approaches have dominated this field, with correlation-based architectures [1, 2] becoming the predominant paradigm. These methods effectively address large-baseline motion by establishing correlation between all pixel pairs using stable appearance features from frames. However, optical flow estimation faces significant challenges despite its success in general scenarios. Frame data becomes unreliable in high dynamic range or rapid motion environments, where most frame-based algorithms fail to perform effectively [3].
Event cameras excel in these challenging scenarios, offering high-frequency motion information that opens new opportunities for visual algorithms [3, 4, 5, 6, 7, 8, 9, 10]. Their unique characteristics facilitate motion estimation across a broader range of scenes, driving advancements in tasks such as motion deblurring [11, 12, 13, 14], object tracking [15, 16], and video interpolation [17, 18, 19, 20, 21]. Recently, event-based optical flow algorithms have gained considerable traction. Unlike frames, event data provides exceptional temporal resolution and dynamic range but lacks consistent visual appearance. State-of-the-art methods [22, 23] leverage the high temporal resolution of events to address issues related to spatial sparsity and noise. Although integrating intermediate motion cues has proven to enhance performance, the inherent sparsity of events often results in unstable feature representation. As illustrated in Fig. 1, event-only methods face challenges such as lack of texture, resulting in structural information loss and inconsistent estimates across consecutive frames. These challenges underscore the need for combining frames and events to improve performance.
@@
To this end, we analyze the advantages of the two modalities in optical flow estimation. Frame features exhibit stable visual appearance in most scenarios, enabling robust spatial correspondences. Event features, on the other hand, provide rich temporal information and retain essential visual structures in challenging environments. Moreover, frames are spatially dense but temporally sparse, whereas events exhibit the opposite characteristics. Building on this observation, we propose leveraging frame data to generate robust spatial correspondences and guide the temporal aggregation of event information. On the other hand, we construct temporally dense correlation maps from event data to extract high temporal resolution motion cues. Our approach integrates the spatial stability of frame data with the temporal richness of event data, enabling effective cross-modal fusion for optical flow estimation.
@@
Figure 1: Improved prediction stability with frame guidance. The top row shows results from the event-only method [22], while the bottom row presents results from our method. These four consecutive frames are taken from the zurich_city_14_c sequence of DSEC-Flow. The lack of texture information in event data results in unstable predictions. Our method leverages stable frame motion features to guide the aggregation of temporally dense event motion features, leading to consistent predictions across consecutive frames.
@@
  • We suggest exploiting the advantages of the spatial robustness of frames and the high temporal resolution of events to improve optical flow estimation.
  • We propose a novel cross-modal collaboration framework that leverages the reconstructed spatially robust modality to complement the sparse event features and guide their temporal aggregation.
  • We produce state-of-the-art performance on the DSEC-Flow benchmark, i.e., compared to the leading fusion-based methods [23], achieving 4% higher accuracy and 45% faster inference.
@@ II-A Learning-based Frame Optical Flow
Figure 2: The overall architecture of our method. The core idea of our approach is to leverage the spatial stability of the frame modality to guide the temporal aggregation of event features, enabling complementary interactions between modalities. To achieve this, an event-enhanced frame representation is introduced as robust spatial guidance. Temporally dense motion cues extracted from events are aggregated into strong motion features, and fed into the iterative process. Furthermore, contextual features from both modalities are fused to enhance optical flow estimation.
@@ II-B Event-based Optical Flow
Event-based optical flow has seen significant advancements over the past decade. Early approaches [31] fit local planes in the event stream to estimate optical flow, while others [32, 33] minimized an energy function using variational optimization. Shiba et al. [34] warps events along point trajectories, deriving an intuitive flow solution. Recently, learning-based methods have demonstrated distinct advantages. Self-supervised methods [35, 36] primarily use contrast maximization [37] or temporal loss [35] to supervise the learning process, while supervised methods rely on labeled data. Notable examples include EvFlowNet [35] and E-RAFT [26], which successfully adapted FlowNet [27] and RAFT [1] architectures for event data. Based on E-RAFT, TMA [22] further utilizes the high temporal resolution of event data to compensate for spatial sparsity. Recently, TCM [38] introduced multi-scale timestamp loss to supervise dense optical flow prediction, achieving notable success. Moreover, Wan et al. [25] and Gehrig et al. [23] integrated image and event data to estimate dense optical flow. Our method emphasizes leveraging the unique characteristics of each modality and their complementary advantages to enable effective cross-modal collaboration.
@@ II-C Event-Frame Cross-Modal Fusion
The complementary nature of events and frames has been extensively explored in various tasks such as semantic segmentation [24, 39, 40, 41], depth estimation [42, 43, 44], and object tracking [45, 46]. In these applications, it is common to employ a multi-stage fusion of event and image features within a UNet-like framework, followed by processing the fused features through the decoder. Chen et al. [46] develop an adaptive sampling method to align event and image modalities, along with a bidirectional-enhanced framework to facilitate cross-modal tracking. Zhu et al. [45] introduce a mask modeling strategy to promote proactive interaction between tokens from different modalities. Several approaches have also explored the integration of events and frames for optical flow estimation. For example, DCEIFlow [25] generates a pseudo-feature for the second frame by fusing the first frame with events and then estimates the optical flow based on the RAFT structure. BFlow [23] concatenates motion features from both images and events into a single representation and then estimates the control point location of the trajectory. Our method guides the aggregation of temporal motion features using spatially robust frame motion features, enabling the fusion of motion cues for optical flow estimation.
@@ III Preliminary and Data Representation
The event and frame data are significantly different and it is necessary to reorganize them. An event is triggered when a pixel of the event camera detects a change in luminance above a threshold $C$. Each event $e_{i}$ typically includes the time $t_{i}$, coordinates $(x_{i},~{}y_{i})$ and the polarity $p_{i}$ of its occurrence. The asynchronous events are usually converted into a frame-like representation. In this paper, we convert the events set into a voxel grid $\mathbf{V}$, as in the previous works [35]:
@@
Different time ranges of events and frame data are required to estimate pixel motion from $T_{k}$ to $T_{k+1}$. For frames data, we use images $I_{k}$ and $I_{k+1}$, corresponding to timestamps $T_{k}$ and $T_{k+1}$. The event data representation is illustrated in Fig. 2. We capture temporally dense motion cues by uniformly partitioning the event stream within the interval $[T_{k},T_{k+1}]$ into a sequence of target segments $\{\mathbf{V}_{1},\mathbf{V}_{2},\ldots,\mathbf{V}_{N}\}$, each with an average time span of $\Delta_{t}$. Then we use the segment from the interval $[T_{k}-\Delta_{t},T_{k}]$ as the reference segment $\mathbf{V}_{0}$. Correlations between the reference and targets provide rich temporal motion information for flow estimation.
@@ IV Proposed Method
Figure 3: Illustration of ICE. The left side shows the event and image reorganization process, with the image captured at time $t$. $\mathbf{V}_{-2}$, $\mathbf{V}_{-1}$, and $\mathbf{V}_{0}$ represent the three voxels proceding $t$. The right side illustrates the overlay of events and frames, demonstrating how their combination enhances spatial robustness.
@@
Specifically, given the input event voxel $\mathbf{V}$ and image $\mathbf{I}$, we first mapped pixel values and voxel values to $[-1,1]$ to balance the contributions of events and frames:
@@
where $\epsilon$ is a small positive constant to prevent the maximum value from being zero. The ICE is then generated as:
@@
where $\mathbf{I}_{t}$ corresponds to time $t$, and $\mathbf{V}_{t-\Delta_{t}}^{t}$ represents the voxel from $t-\Delta_{t}$ to $t$.
As shown in Fig. 3, ICE combines intensity information from images with structural information from events. In low-light conditions, frame data may degrade significantly, while event data still provide sufficient spatial information. In well-lit scenes, frame intensity information complements the event data. Consequently, the motion features extracted from ICE exhibit high spatial stability, ensuring robust guidance. Note that we use events only within a time window of length $\Delta t$ to avoid edge blurring caused by long windows.
@@
where feature $\mathbf{F}_{1}$ corresponding to the first ICE, and $D$ is the feature channels. We emphasize that $\mathbf{C}_{g}\in\mathbb{R}^{HW\times HW}$ encompasses not only the motion information for the entire optical flow from $T_{k}$ to $T_{k+1}$ but also the robust spatial correlation generated from the ICE, thereby supporting the guidance for event features.
Figure 4: The architecture of the Mix Fusion Context Encoder. We extract temporal and spatial contextual features from events and images, respectively, and fuse them into spatiotemporal features. Spatial context is extracted from the first frame $I_{k}$, while temporal context includes all events from $T_{k}$ to $T_{k+1}$.
Temporally Dense Correlation. The effectiveness of leveraging the high temporal resolution of event data to compensate for extreme sparsity and noise has been validated in several works. TMA [22] and BFlow [23] propose to replace the single correlation volume with temporally-dense correlation volumes and achieved excellent accuracy. Thus we extract event motion cues at multiple intermediate times, following these works.
Specifically, we first split the event stream into a reference $\mathbf{V}_{0}$ and a series of targets $\{\mathbf{V}_{n},n\in[1,N]\}$, as described in Sec. III. The reference $\mathbf{V}_{0}$ and the final target $\mathbf{V}_{N}$ are aligned with the first and second frames $I_{k}$ and $I_{k+1}$, respectively. To ensure consistency, all the targets and the reference share the same feature encoder. The correlation volumes are then generated between the reference and each target, to construct the temporally dense cost volumes:
@@ IV-B Context Fusion and Guided Aggregation
To achieve this, we propose the Mix-Fusion Context Encoder, illustrated in Fig. 4, to extract spatiotemporal context features from events and frames. We first extract spatial context features from $I_{k}$ and temporal context features from $\mathbf{V}_{T_{k}}^{T_{k+1}}$ using two separate encoders. Both features are designed to have the same channel dimensions and are subsequently fed into the mix-fusion block. The mix-fusion module is inspired by [47], and the output spatiotemporal feature $\mathbf{F}_{st}$ has the same number of channels as single-modal context features:
@@
Figure 5: Illustration of the transformer-based cross-modal guided aggregation strategy. The robust ICE feature (blue) is used as the Key to guide the aggregation of temporally dense event features (green) and also serves as the Value to enhance the spatially sparse event features.
@@
$Cost=\texttt{lookup}(\mathbf{C}_{g},\hat{f}),$
where $\texttt{lookup}(\cdot,\cdot)$ samples the similarity between pixel pairs from $\mathbf{C}_{g}$ based on $\hat{f}$. For the events modality, each reference-target pair has different time intervals, necessitating a linear lookup strategy [22]. Given the flow estimation $\hat{f}$, the $i$-th cost map is sampled as:
@@
The sampled cost maps are encoded as motion features by $E_{M}$, as shown in Fig. 2. The ICE motion feature $\mathbf{M}_{ice}$ encodes robust visual information, while the event motion features $\{\mathbf{M}_{ev}^{i},i\in[1,N]\}$ provide temporally detailed motion cues. Subsequently, we use $\mathbf{M}_{ice}$ to guide the aggregation of $\mathbf{M}_{ev}$. Inspired by [2, 22], we adopt a transformer to implement the cross-modal guided aggregation:
@@
where $D$ is the dimension of $\mathbf{K}$, and $\texttt{ffn}(\cdot)$ is the feed-forward neural network. The $\mathbf{M}_{ice}$ functions as both the Key and the Value, guiding the aggregation of temporal event features and compensating for spatial sparsity in the event data. Eventually, the aggregated features are concatenated into a single motion feature, which is fed into the ConGRU for further refinement.

### v2

@@ Abstract
Current optical flow methods exploit the stable appearance of frame (or RGB) data to establish robust correspondences across time. Event cameras, on the other hand, provide high-temporal-resolution motion cues and excel in challenging scenarios. These complementary characteristics underscore the potential of integrating frame and event data for optical flow estimation. However, most cross-modal approaches fail to fully utilize the complementary advantages, relying instead on simply stacking information. This study introduces a novel approach that uses a spatially dense modality to guide the aggregation of the temporally dense event modality, achieving effective cross-modal fusion. Specifically, we propose an event-enhanced frame representation that preserves the rich texture of frames and the basic structure of events. We use the enhanced representation as the guiding modality and employ events to capture temporally dense motion information. The robust motion features derived from the guiding modality direct the aggregation of motion information from events. To further enhance fusion, we propose a transformer-based module that complements sparse event motion features with spatially rich frame information and enhances global information propagation. Additionally, a mix-fusion encoder is designed to extract comprehensive spatiotemporal contextual features from both modalities. Extensive experiments on the MVSEC and DSEC-Flow datasets demonstrate the effectiveness of our framework. Leveraging the complementary strengths of frames and events, our method achieves leading performance on the DSEC-Flow dataset. Compared to the event-only model, frame guidance improves accuracy by 10%. Furthermore, it outperforms the state-of-the-art fusion-based method with a 4% accuracy gain and a 45% reduction in inference time. The code is publicly available at https://github.com/ZhouQianang/STFlow.
@@ I Introduction
Optical flow estimation plays a vital role in understanding object motion between image pairs, providing valuable insights into scene dynamics. Frame-based optical flow methods have leveraged the consistent visual appearance of images, leading to extensive development. Over the past decade, learning-based approaches have dominated this field, with correlation-based architectures [1, 2] becoming the predominant paradigm. These methods effectively address large-baseline motion by establishing the correlation between all pixel pairs using stable appearance features from frames. However, optical flow estimation faces significant challenges despite its success in general scenarios. Frame data becomes unreliable in high dynamic range or rapid motion environments, where most frame-based algorithms fail to perform effectively [3].
Figure 1: Frame-guided prediction improves stability. The lack of texture in event data leads to unstable predictions. With the proposed strategy that guides the aggregation of temporal features using spatially stable frame, our method achieves more consistent predictions across frames.
Event cameras excel in these challenging scenarios, offering high-frequency motion information that opens new opportunities for visual algorithms [3, 4, 5, 6, 7, 8, 9, 10]. Their unique characteristics facilitate motion estimation across a broader range of scenes, driving advancements in tasks such as motion deblurring [11, 12, 13, 14], object tracking [15, 16], and video interpolation [17, 18, 19, 20, 21]. Recently, event-based optical flow algorithms have gained considerable traction. Unlike frames, event data provides exceptional temporal resolution and dynamic range but lacks a consistent visual appearance. State-of-the-art methods [22, 23] leverage the high temporal resolution of events to address issues related to spatial sparsity and noise. Although integrating intermediate motion cues has proven to enhance performance, the inherent sparsity of events often results in unstable feature representation. As illustrated in Fig. 1, event-only methods face challenges such as a lack of texture, resulting in structural information loss and inconsistent estimates across consecutive frames. These challenges underscore the need for combining frames and events to improve performance.
@@
To this end, we analyze the advantages of the two modalities in optical flow estimation. Frame features exhibit a stable visual appearance in most scenarios, enabling robust spatial correspondences. Event features, on the other hand, provide rich temporal information and retain essential visual structures in challenging environments. Moreover, frames are spatially dense but temporally sparse, whereas events exhibit the opposite characteristics. Building on this observation, we propose leveraging frame data to generate robust spatial correspondences and guide the temporal aggregation of event information. On the other hand, we construct temporally dense correlation maps from event data to extract high temporal resolution motion cues. Our approach integrates the spatial stability of frame data with the temporal richness of event data, enabling effective cross-modal fusion for optical flow estimation.
@@
  • We propose a cross-modal collaboration framework that performs multi-level, modality-specific fusion of frame and event, effectively leveraging their complementary characteristics to enhance optical flow estimation.
  • We design a lightweight spatially robust modality and a spatially guided temporal aggregation strategy to facilitate high-level feature fusion between frame and event.
  • Our method achieves state-of-the-art performance on the DSEC-Flow benchmark and remains competitive with existing supervised approaches on the MVSEC dataset.
@@ II-A Learning-based Frame Optical Flow
Figure 2: The overall architecture of our method. The core idea of our approach is to leverage the spatial stability of the frame modality to guide the temporal aggregation of event features, enabling complementary interactions between modalities. To achieve this, an event-enhanced frame representation ICE is introduced as robust spatial guidance. The ICE feature is then used as the $K$ to guide the aggregation of temporally dense event features and also serves as the $V$ to enhance the spatially sparse events. Furthermore, contextual features from both modalities are fused to enhance optical flow estimation.
@@ II-B Event-based Optical Flow
Event-based optical flow has seen significant advancements in the past decade. Early approaches [31] fit local planes in the event stream to estimate optical flow, while others [32, 33] minimized an energy function using variational optimization. Shiba et al. [34] warps events along point trajectories, deriving an intuitive flow solution. They further extended this optimization framework to a broader range of event-based vision tasks in [35]. Recently, learning-based methods have demonstrated distinct advantages. Self-supervised methods [36, 37] primarily use contrast maximization [38] or temporal loss [36] to supervise the learning process, while supervised methods rely on labeled data. Notable examples include EvFlowNet [36] and E-RAFT [26], which successfully adapted FlowNet [27] and RAFT [1] architectures for event data. Based on E-RAFT, TMA [22] further utilizes the high temporal resolution of event data to compensate for spatial sparsity. Recently, TCM [39] introduced multi-scale timestamp loss to supervise dense optical flow prediction, achieving notable success. Moreover, Wan et al. [25] and Gehrig et al. [23] integrated image and event data to estimate dense optical flow. Our method emphasizes the use of the unique characteristics of each modality and their complementary advantages to enable effective cross-modal collaboration.
@@ II-C Event-Frame Cross-Modal Fusion
The complementary nature of events and frames has been extensively explored in various tasks such as semantic segmentation [24, 40, 41, 42], depth estimation [43, 44, 45], and object tracking [46, 47]. In these applications, it is common to employ a multi-stage fusion of event and image features within a UNet-like framework, followed by processing the fused features through the decoder. Chen et al. [47] develop an adaptive sampling method to align event and image modalities, along with a bidirectional-enhanced framework to facilitate cross-modal tracking. Zhu et al. [46] introduce a mask modeling strategy to promote proactive interaction between tokens from different modalities. Several approaches have also explored the integration of events and frames for optical flow estimation. For example, DCEIFlow [25] generates a pseudo-feature for the second frame by fusing the first frame with events and then estimates the optical flow based on the RAFT structure. BFlow [23] concatenates motion features from both images and events into a single representation and then estimates the control point location of the trajectory. Our method guides the aggregation of temporal motion features using spatially robust frame motion features, enabling the fusion of motion cues.
@@ III Preliminary and Data Representation
The event and frame data are significantly different, and it is necessary to reorganize them. An event is triggered when a pixel of the event camera detects a change in luminance above a threshold $C$. Each event $e_{i}$ typically includes the time $t_{i}$, coordinates $(x_{i},~y_{i})$ and the polarity $p_{i}$ of its occurrence. The asynchronous events are usually converted into a frame-like representation. In this paper, we convert the event set into a voxel grid $\mathbf{V}$, as in the previous works [36]:
@@
Different time ranges of events and frame data are required to estimate pixel motion from $T_{k}$ to $T_{k+1}$. For frames data, we use images $I_{k}$ and $I_{k+1}$, corresponding to timestamps $T_{k}$ and $T_{k+1}$. The event data representation is illustrated in Fig. 2. We capture temporally dense motion cues by uniformly partitioning the event stream within the interval $T_{k},T_{k+1}]$ into a sequence of target segments $\{\mathbf{V}_{1},\mathbf{V}_{2},\ldots,\mathbf{V}_{N}\}$, each with an average time span of $\Delta t$. Then we use the segment from the interval $T_{k}-\Delta t,T_{k}]$ as the reference segment $\mathbf{V}_{0}$. Correlations between the reference and targets provide rich temporal motion information for flow estimation.
@@ IV Proposed Method
Figure 3: Complementary strengths of frames and events across scenes. Frames provide rich texture under normal lighting but often fail in challenging illumination, where events preserve clear structural information.
@@
Specifically, given the input event voxel $\mathbf{V}$ and image $\mathbf{I}$, we first mapped pixel values and voxel values to [1,1][-1,1] to balance the contributions of events and frames:
@@
where $\epsilon$ is a small positive constant to prevent the maximum value from being zero, and $\max(|\mathbf{V}|)$ is computed individually for each sample. The ICE is then generated as:
@@
where $\mathbf{I}_{t}$ corresponds to time $t$, and $\mathbf{V}_{t-\Delta t}^{t}$ represents the voxel from $t-\Delta t$ to $t$.
As shown in Fig. 3, events and frames exhibit strong complementarity. In low-light conditions, frame data may degrade significantly, while event data still provides sufficient spatial information. In well-lit scenes, frame intensity information complements the event data. Consequently, the motion features extracted from ICE exhibit high spatial stability, ensuring robust guidance. Note that we use event only within a time window of length $\Delta t$ to avoid edge blurring caused by long windows. The event time window of ICE matches the temporal length in the event target. This design maintains consistency during cross-modal aggregation while avoiding unnecessary computational overhead.
@@
where $\mathbf{F}_{1},\mathbf{F}_{2}$ denote the features corresponding to the start and end ICEs, and $D$ is the feature channels. We emphasize that $\mathbf{C}_{G}\in\mathbb{R}^{HW\times HW}$ encompasses not only the motion information for the entire optical flow from $T_{k}$ to $T_{k+1}$ but also the robust spatial correlation generated from the ICE, thereby supporting the guidance for event features.
Temporally Dense Correlation. The effectiveness of leveraging the high temporal resolution of event data to compensate for sparsity and noise has been validated in several recent works. Notably, TMA [22] and BFlow [23] improve performance by replacing the single correlation volume with temporally-dense correlation volumes. We believe this design benefits from the temporal consistency of structural motion patterns, while event noise, especially in low-light conditions, is generally uncorrelated across time. Aggregating correlation information across multiple time steps thus enhances the signal-to-noise ratio by reinforcing consistent features and suppressing transient noise. Building on this insight, we extract event motion cues at multiple intermediate times, following these works.
Specifically, we first split the event stream into a reference $\mathbf{V}_{0}$ and a series of targets $\{\mathbf{V}_{n},n\in[1,N]\}$, as described in Sec. III. The reference $\mathbf{V}_{0}$ and the final target $\mathbf{V}_{N}$ are aligned with the first and second frames $I_{k}$ and $I_{k+1}$, respectively. To ensure consistency, all the targets and the reference share the same feature encoder. The correlation volumes are then generated between the reference and each target to construct the temporally dense cost volumes:
@@ IV-B Context Fusion and Guided Aggregation
To achieve this, we propose the Mix-Fusion Context Encoder, illustrated in Fig. 2, to extract spatiotemporal context features from events and frames. We first extract spatial context features from $I_{k}$ and temporal context features from the event voxel between $T_{k}$ and $T_{k+1}$, using two separate encoders. Both features are designed to have the same channel dimensions and are subsequently fed into the mix-fusion block. The mix-fusion module is inspired by [48], and the output spatiotemporal feature $\mathbf{F}_{st}$ has the same number of channels as single-modal context features:
@@
[…]
where $\texttt{lookup}(\cdot,\cdot)$ samples the similarity between pixel pairs from $\mathbf{C}_{G}$ based on $\hat{f}$. For the event modality, each reference-target pair has different time intervals, necessitating a linear lookup strategy [22]. Given the flow estimation $\hat{f}$, the $i$-th cost map is sampled as:
@@
The sampled cost maps are encoded as motion features by $E_{M}$, as shown in Fig. 2. The ICE motion feature $\mathbf{M}_{ice}$ encodes robust visual information, while the event motion features $\{\mathbf{M}_{ev}^{i},i\in[1,N]\}$ provide temporally detailed motion cues. Subsequently, we use $\mathbf{M}_{ice}$ to guide the aggregation of $\mathbf{M}_{ev}$. Inspired by [2, 22], we adopt a lightweight Transformer to implement the cross-modal guided aggregation. Event motion features $\mathbf{M}_{ev}$ are used as queries, while ICE motion features $\mathbf{M}_{ice}$ serve as both keys and values, guiding and facilitating the event features through token-level cross-attention:
@@
where $D$ is the dimension of $\mathbf{K}$, and $\texttt{ffn}(\cdot)$ is the feed-forward neural network. In our experiments, we use a single-head, single-layer configuration with shared embedding dimensions. Eventually, the aggregated features are concatenated into a single motion feature, which is fed into the ConGRU for further refinement.